# Physics of droplet regulation in biological cells

### v1

@@ Abstract
Droplet formation has emerged as an essential concept for the spatiotemporal organisation of biomolecules in cells. However, classical descriptions of droplet dynamics based on passive liquid-liquid phase separation cannot capture the complex situations inside cells. This review discusses three general aspects that are crucial in cells: (i) biomolecules are diverse and individually complex, implying that cellular droplets posses complex internal behaviour, e.g., in terms of their material properties; (ii) the cellular environment contains many solid-like structures that droplets can wet; (iii) cells are alive and use fuel to drive processes out of equilibrium. We illustrate how these principles control droplet nucleation, growth, position, and count to unveil possible regulatory mechanisms in biological cells and other applications of phase separation.
  • Version of draft: January 23, 2025
@@ 1 Introduction
The physics of droplet formation, known as phase separation, has been studied for more than a century [19], including in biological contexts [20]. Phase separation generally describes how a balance between entropy, favouring mixing, and enthalpy, due to interactions, leads to stable droplets in thermodynamic equilibrium. While this basic principle likely also governs droplet formation in cells [21], phase separation in a biological context is significantly more complex for at least three major reasons (figure 1.1). First, cells comprise an enormous number of different biomolecules, which can each be incredibly complex [22]. Consequently, droplets typically comprise many components and posses complex material properties. Second, the cellular environment is heterogeneous, filled with membranes, rigid polymers, and other macromolecular structures. Droplets interact with this environment through wetting, which affects both the droplets and the environment itself. Third, cells are alive, implying that processes can be driven away from equilibrium, which can profoundly affect droplet formation [23]. One challenge in describing cellular droplets theoretically lies in the many different length scales involved: microscopic interactions (e.g., $\pi$-$\pi$-stacking, cation-$\pi$ interactions, hydrogen bonding, and hydrophobic interactions [24]) induce phase separation; mesoscopic interactions (e.g., complex coacervation [25], entanglement, and gelation) determine mechanical properties of droplets; and macroscopic interactions (e.g., gradients, flow, and wetting) control droplets on the cellular scale. Theory is necessary to bridge these scales and to arrive at a holistic description of the physical processes shaping droplets in cells.
@@
This review is organised into four distinct sections, beginning with a basic summary of the physics of phase separation in section 2. The three subsequent sections are intended to be independent of each other and can be read in any order. These sections cover the core aspects of cellular complexity identified above (figure 1.1): internal complexity of droplets (section 3), complexity of the surrounding environment (section 4), and activity (section 5).
@@ 2.1.1  Thermodynamics of liquids
We first consider an isothermal, homogeneous system of volume $V_{\mathrm{sys}}$ filled with a liquid comprised of a solvent and ${N_{\mathrm{c}}}$ additional components. For simplicity, we focus here on incompressible systems, where the molecular volumes $\nu_{i}$ of each component are constant. The generalisation to compressible systems is discussed in appendix A. The composition of the homogeneous system is fully described by the particle counts $N_{i}$ for $i=1,\ldots,{N_{\mathrm{c}}}$ and all remaining space is simply filled with solvent particles. Since the system is isothermal and its volume is constant, the equilibrium thermodynamics follow from the Helmholtz free energy $F=V_{\mathrm{sys}}f(\phi_{1},\ldots,\phi_{N_{\mathrm{c}}})$, where $f$ is the free energy density which depends on the ${N_{\mathrm{c}}}$ volume fractions $\phi_{i}=\nu_{i}N_{i}/V_{\mathrm{sys}}$. We consider the Flory–Huggins free energy [94, 95],
@@
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$ due to incompressibility; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, which can be determined from detailed theories [81]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and solvent, but it can also account for external potentials. These potentials can in principle depend on space, like gravity or regulator gradients [23], but we do not discuss this further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A.
To study thermodynamic equilibria, we derive intensive thermodynamic variables using derivatives of the free energy $F$ with respect to the extensive variables $N_{i}$ and $V_{\mathrm{sys}}$. Note that these variables are not independent, since $V_{\mathrm{sys}}=\sum_{i=0}^{N_{\mathrm{c}}}\nu_{i}N_{i}$ in the incompressible limit. Consequently, the chemical potentials $\bar{\mu}_{i}=(\partial F/\partial N_{i})_{N_{j\neq i},V_{\mathrm{sys}}}$,
@@
are actually exchange chemical potentials, quantifying the change of free energy when $\Omega_{i}$ solvent particles are replaced by a single particle of type $i$. Similarly, the pressure $\Pi=-(\partial F/\partial V_{\mathrm{sys}})_{N_{i}}=\sum_{i}\bar{\mu}_{i}\phi_{i}\nu_{i}^{-1}-f$,
@@
$\bar{\mu}_{i}\bigl{(}\Phi^{(1)}\bigr{)}$ $=\bar{\mu}_{i}\bigl{(}\Phi^{(2)}\bigr{)}=\cdots=\bar{\mu}_{i}\bigl{(}\Phi^{({N_{\mathrm{p}}})}\bigr{)}$ (2.5a)
@@
are met for components $i=1,\ldots,{N_{\mathrm{c}}}$. These equations denote chemical and mechanical equilibrium between phases, respectively. These are $({N_{\mathrm{p}}}-1)({N_{\mathrm{c}}}+1)$ conditions for ${N_{\mathrm{p}}}{N_{\mathrm{c}}}$ independent fractions $\phi^{(n)}_{i}$. The system is thus overdetermined when ${N_{\mathrm{p}}}>{N_{\mathrm{c}}}+1$, illustrating that at most ${N_{\mathrm{c}}}+1$ phases of different composition can coexist, consistent with Gibbs phase rule for fixed temperature and system volume [19].
@@ 2.1.2  Kinetics of heterogeneous systems
To describe droplet dynamics, we need to generalise the theory of the previous section to heterogeneous systems. To do this, we assume local equilibrium, neglecting heat fluctuations which manifest on long length and time scales in these typically active systems [98]. We can then assume an isothermal system and define fields of the thermodynamic quantities introduced above. In particular, the system’s state is now specified by volume fraction fields $\Phi(\boldsymbol{r})=\bigl{(}\phi_{1}(\boldsymbol{r}),\ldots,\phi_{{N_{\mathrm{c}}}}(\boldsymbol{r})\bigr{)}$ and its free energy is given by the functional [99, 100, 101]
@@
We derive the kinetics based on linear non-equilibrium thermodynamics [96], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ (assuming equal mass density for all components) and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Hence,
@@
for $i=1,\ldots,{N_{\mathrm{c}}}$, and the solvent flux is given by $\boldsymbol{j}_{0}=-\sum_{i=1}^{N_{\mathrm{c}}}\boldsymbol{j}_{i}$, so that all relative fluxes add to zero. The thermal fluctuations described by $\boldsymbol{\xi}_{i}$ have zero mean and their correlations are governed by the fluctuation-dissipation theorem [96],
@@
where $\mathbb{1}$ denotes the identity matrix capturing that spatial components of $\boldsymbol{\xi}_{i}$ are independent. Onsager’s symmetry principle implies $\Lambda_{ij}=\Lambda_{ji}$, but to derive expressions for the ${N_{\mathrm{c}}}$-dimensional matrix $\Lambda_{ij}$ it is instructive to express it by the more general $({N_{\mathrm{c}}}+1)$-dimensional mobility matrix $L_{ij}$, which captures the interplay of all components. We show in appendix A that the two matrices are connected by
@@
The two matrices are identical in the simple case where the solvent is fast ($L_{00}\rightarrow\infty$). In contrast, in the case where all components have equal molecular sizes ($\Omega_{i}=1$) and mobilities ($L_{ij}=\lambda\phi_{i}\delta_{ij}$), equation (2.12) implies effective cross-diffusion [104, 105, 103],
@@
which captures the effects of the kinetics of the solvent component, which is not described explicitly. More realistic kinetic models capture additional concentration-dependences [105, 106, 107]. In all these cases, the thermal fluctuations described by equation (2.11) are then correlated, which can be simulated using Cholesky decomposition [108]. Taken together, we find
@@
which is augmented by equation (2.9) for the flow field $\boldsymbol{v}$. However, flows can be neglected ($\boldsymbol{v}=\boldsymbol{0}$) when droplet growth is dominated by diffusion. For example, C. elegans centrosomes and P granules of typical radii $R\approx 1\,\mathrm{\upmu m}$ [109, 110] exhibit velocities of $v\approx 10\,\nicefrac{\mathrm{\upmu m}}{\mathrm{min}}$ [111, 112], but the molecular diffusivity of $D\approx 5\,\nicefrac{\mathrm{\upmu m^{2}}}{\mathrm{s}}$ [113, 114] implies small Péclet numbers ($vR/D\approx 0.03$), so hydrodynamic transport is less relevant. Consequently, we also neglect the stochastic component of the hydrodynamic flow [115].
The coupled partial differential equations (2.14) for the fields $\{\phi_{i}\}$ typically permit stationary homogeneous states. For instance, without chemical reactions ($s_{i}=0$), each homogeneous state $\phi_{i}(\boldsymbol{r})=\bar{\phi}_{i}$ is stationary because of material conservation. Such a state is stable if the associated Hessian
@@
The late-stage dynamics can in principle be obtained by simulating equations (2.14), which requires two boundary conditions at the system’s boundary because of the fourth-order derivatives. Before we introduce such boundary conditions in section 4, we will consider the simpler case of isolated droplets, where we can employ periodic boundary conditions on all fields. This situation is typically considered in numerical investigations, e.g., using finite difference [116], spectral methods [117, 118], Lattice-Boltzmann methods [119, 120, 121, 122], or stochastic solvers [123, 124]. Since these numerical simulations are costly [125], contemporary computers are not able to simulate multi-component dynamics for the time and length scales relevant in cells [116]. However, to discuss dynamics of individual droplets, it often suffices to consider binary phase separation and sharp interfaces (see section 2.2), which is the main focus of this review.
@@ 2.1.3  Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black lines) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) for two values of the Laplace pressure $P_{\gamma}$. The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium concentrations (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
In many cases, we do not need to describe all the individual components of a mixture and instead focus on droplet material separating from a solvent. In this case of a binary liquid, only the fraction $\phi$ of one component is specified (${N_{\mathrm{c}}}=1$) and the free energy density reduces to
@@
and determines the phase behaviour of the liquid; see figure 2.1. In particular, a homogeneous system with fraction $\phi$ is stable with respect to small perturbations when the Hessian given in equation (2.15) is negative, i.e., if $f^{\prime\prime}(\phi)\propto\bar{\mu}^{\prime}(\phi)>0$. This occurs when the interaction parameter $\chi$ is smaller than
@@
Homogeneous systems are stable against small fluctuations outside the spinodal region, but they are not necessarily the only stable states or even the state with the lowest free energy. To see whether heterogeneous states involving droplets can form, we next consider the equilibrium conditions (2.5). These conditions can be solved graphically using the common-tangent construction shown in figure 2.1A, which is equivalent to the equal-area Maxwell construction shown in figure 2.1B. Alternatively, the equilibrium point can be read off the $\Pi(\bar{\mu})$ plot shown in figure 2.1D, and the corresponding volume fractions then follow from figure 2.1B–C. For the chemical potential of the binary system given by equation (2.17), these graphical constructions work, and the solutions are unique, if $\chi>\chi_{*}$. The two fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ correspond to the equilibrium fractions of the dilute and dense phase, respectively outside and inside droplets. Note that $\phi^{(0)}_{\mathrm{out}}$ is also known as the saturation concentration since droplets can only form when it is exceeded. The common-tangent construction also illustrates that the phase-separated state generally has lower free energy compared to the homogeneous state at equal average composition $\bar{\phi}$. The two states only have equal free energy in the limit where one phase of the phase separated state vanishes (when $\bar{\phi}=\phi^{(0)}_{\mathrm{out}}$ or $\bar{\phi}=\phi^{(0)}_{\mathrm{in}}$). Taken together, the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ as a function of $\chi$ form the binodal line. Since the phase-separated state does not exist outside the binodal line, this line also bounds the region where the homogeneous phase is globally stable; see figure 2.1. The binodal line can be obtained analytically for the symmetric case,
@@
for $\Omega=1$. In the limit of strong interactions ($\chi\gg 1$), this predicts $\phi^{(0)}_{\mathrm{out}}\approx e^{-\chi}$, which reveals that even a minute increase of the interaction energy by $2k_{\mathrm{B}}T$ (so $\chi$ is increased by $2$) leads to almost a $10$-fold decrease in the saturation concentration $\phi^{(0)}_{\mathrm{out}}$. In contrast, the spinodal branch scales as $\phi^{\mathrm{spin}}_{\mathrm{out}}\approx 1/(2\Omega\chi)$, implying that a linear stability analysis of the homogeneous state is inadequate to study phase separation in detail.
The entire phase diagram can be separated into three phases: for low $\chi$ or dilute solutions ($\phi<\phi^{(0)}_{\mathrm{out}}$ or $\phi>\phi^{(0)}_{\mathrm{in}}$) only homogeneous states persist. Conversely, only the phase separate state exists in the spinodal region when interactions are sufficiently strong ($\chi>\chi_{\mathrm{spin}}$). Between these two regions, homogeneous and phase separated states are both locally stable, but the phase separated state has lower energy; nucleation events become possible when thermal fluctuations are included.
@@
where we for simplicity neglect hydrodynamic flows $\boldsymbol{v}$, and the mobility $\Lambda(\phi)=\Lambda_{0}\phi(1-\phi)$ follows from equation (2.13). Without the source term $s$, accounting for reactions described in section 5, equation (2.23) is know as the Cahn–Hilliard equation [99]. This equation describes many phase separation phenomena and has been studied extensively [127].

### v2

@@ Abstract
Droplet formation has emerged as an essential concept for the spatiotemporal organisation of biomolecules in cells. However, classical descriptions of droplet dynamics based on passive liquid-liquid phase separation cannot capture the complex situation inside cells. This review discusses three distinct aspects that are crucial in cells: (i) biomolecules are diverse and individually complex, implying that cellular droplets posses complex internal behaviour, e.g., in terms of their material properties; (ii) the cellular environment contains many solid-like structures that droplets can wet; (iii) cells are alive and use fuel to drive processes out of equilibrium. We illustrate how these principles control droplet nucleation, growth, position, and count to unveil possible regulatory mechanisms in biological cells and other applications of phase separation.
  • Version of draft: February 27, 2025
@@ 1 Introduction
The physics of droplet formation, known as phase separation, has been studied for more than a century [19], including in biological contexts [20]. Phase separation generally describes how a balance between entropy, favouring mixing, and enthalpy, due to interactions, leads to stable droplets in thermodynamic equilibrium. While this basic principle likely also governs droplet formation in cells [21], phase separation in a biological context is significantly more complex for at least three major reasons (figure 1.1). First, cells comprise an enormous number of different biomolecules, which can each be incredibly complex [22]. Consequently, droplets typically comprise many components and posses complex material properties. Second, the cellular environment is heterogeneous, filled with membranes, rigid polymers, and other macromolecular structures. Droplets interact with this environment through wetting, which affects both the droplets and the environment itself. Third, cells are alive, implying that processes can be driven away from equilibrium, which can profoundly affect droplet formation [23]. One challenge in describing cellular droplets theoretically lies in the many different length scales involved: microscopic interactions (e.g., $\pi$-$\pi$-stacking, cation-$\pi$ interactions, hydrogen bonding, and hydrophobic interactions [24]) induce phase separation; mesoscopic interactions (e.g., complex coacervation [25], entanglement, and gelation) determine mechanical properties of droplets; and macroscopic interactions (e.g., gradients, flow, and wetting) control droplets on the cellular scale. Theory is necessary to bridge these scales and to arrive at a holistic description of the physical processes that shape droplets in cells.
@@
This review is organised into four distinct sections, beginning with a summary of the basic physics of phase separation in section 2. The three subsequent sections are intended to be independent of each other and can be read in any order. These sections cover the core aspects of cellular complexity identified above (figure 1.1): internal complexity of droplets (section 3), complexity of the surrounding environment (section 4), and activity (section 5).
@@ 2.1.1  Thermodynamics of liquids
We first consider an isothermal, homogeneous system of volume $V_{\mathrm{sys}}$ filled with a liquid comprised of a solvent and ${N_{\mathrm{c}}}$ additional components. For simplicity, we focus here on incompressible systems, where the molecular volumes $\nu_{i}$ of each component are constant. The generalisation to compressible systems is discussed in appendix A. The composition of the homogeneous system is fully described by the particle counts $N_{i}$ for $i=1,\ldots,{N_{\mathrm{c}}}$ and all remaining space is filled with solvent. Since the system is isothermal and has constant volume, the relevant thermodynamic potential is the Helmholtz free energy $F=V_{\mathrm{sys}}f(\phi_{1},\ldots,\phi_{N_{\mathrm{c}}})$, where $f$ is the free energy density which is a function of the ${N_{\mathrm{c}}}$ volume fractions $\phi_{i}=\nu_{i}N_{i}/V_{\mathrm{sys}}$. We consider here the Flory–Huggins free energy [94, 95],
@@
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$ due to incompressibility; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms in equation (2.1) specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, whose particular forms can be determined from detailed theories [81]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and the solvent, but it can also account for external potentials. Such potentials can in principle depend on space, due to, for example, gravity or gradients in a regulatory species [23], but we do not discuss this aspect further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A.
To study thermodynamic equilibria, we derive intensive thermodynamic variables using derivatives of the free energy $F$ with respect to the extensive variables $N_{i}$ and $V_{\mathrm{sys}}$. Note that these variables are not independent, since $V_{\mathrm{sys}}=\sum_{i=0}^{N_{\mathrm{c}}}\nu_{i}N_{i}$ in the incompressible limit. Consequently, the relevant chemical potentials $\bar{\mu}_{i}=(\partial F/\partial N_{i})_{N_{j\neq i},V_{\mathrm{sys}}}$,
@@
are exchange chemical potentials, quantifying the change of free energy when $\Omega_{i}$ solvent particles are replaced by a single particle of type $i$. Similarly, the pressure $\Pi=-(\partial F/\partial V_{\mathrm{sys}})_{N_{i}}=\sum_{i}\bar{\mu}_{i}\phi_{i}\nu_{i}^{-1}-f$,
@@
$\bar{\mu}_{i}\bigl{(}\Phi^{(1)}\bigr{)}$ $=\bar{\mu}_{i}\bigl{(}\Phi^{(2)}\bigr{)}=\cdots=\bar{\mu}_{i}\bigl{(}\Phi^{({N_{\mathrm{p}}})}\bigr{)}\;,$ (2.5a)
@@
are met for components $i=1,\ldots,{N_{\mathrm{c}}}$. These equations represent chemical and mechanical equilibrium between phases, respectively. Equations (2.5) give $({N_{\mathrm{p}}}-1)({N_{\mathrm{c}}}+1)$ conditions for ${N_{\mathrm{p}}}{N_{\mathrm{c}}}$ independent fractions $\phi^{(n)}_{i}$. The system is thus overdetermined when ${N_{\mathrm{p}}}>{N_{\mathrm{c}}}+1$, illustrating that at most ${N_{\mathrm{c}}}+1$ phases of different composition can coexist, consistent with Gibbs phase rule for fixed temperature and system volume [19].
@@ 2.1.2  Kinetics of heterogeneous systems
To describe droplet dynamics, we must generalise the theory presented in the previous section to heterogeneous systems. To do this, we assume local equilibrium, neglecting the heat fluctuations which can manifest on long length and time scales in these typically active systems [98]. We can thus assume an isothermal system and define fields for the thermodynamic quantities introduced above. In particular, the system state is now specified by volume fraction fields $\Phi(\boldsymbol{r})=\bigl{(}\phi_{1}(\boldsymbol{r}),\ldots,\phi_{{N_{\mathrm{c}}}}(\boldsymbol{r})\bigr{)}$ and its free energy is given by the functional [99, 100, 101]
@@
We now derive kinetic equations based on linear non-equilibrium thermodynamics [96], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions as described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ (assuming equal mass density for all components) and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Hence,
@@
for $i=1,\ldots,{N_{\mathrm{c}}}$, and the solvent flux is given by $\boldsymbol{j}_{0}=-\sum_{i=1}^{N_{\mathrm{c}}}\boldsymbol{j}_{i}$, so that all relative fluxes sum to zero. The thermal fluctuations described by $\boldsymbol{\xi}_{i}$ have zero mean and their correlations are governed by the fluctuation–dissipation theorem [96],
@@
where $\mathbb{1}$ denotes the identity matrix, indicating that spatial components of $\boldsymbol{\xi}_{i}$ are independent. Onsager’s symmetry principle implies $\Lambda_{ij}=\Lambda_{ji}$, but to derive expressions for the ${N_{\mathrm{c}}}$-dimensional matrix $\Lambda_{ij}$ it is instructive to express it by the more general $({N_{\mathrm{c}}}+1)$-dimensional mobility matrix $L_{ij}$, which captures the interplay of all components. We show in appendix A that the two matrices are connected by
@@
The rescaled mobility matrix $\Lambda_{ij}$ is identical to $L_{ij}$ in the simple case where the solvent is fast ($L_{00}\rightarrow\infty$). In contrast, even in the simple case where all components have equal molecular sizes ($\Omega_{i}=1$) and mobilities scale with density ($L_{ij}=\lambda\phi_{i}\delta_{ij}$), equation (2.12) implies effective cross-diffusion [104, 105, 103],
@@
which captures the effects of the kinetics of the solvent component not described explicitly. More realistic kinetic models capture additional concentration-dependences [105, 106, 107]. In all these cases, the thermal fluctuations described by equation (2.11) are then correlated, which can be simulated using Cholesky decomposition [108]. Taken together, we find
@@
which is augmented by equation (2.9) for the flow field $\boldsymbol{v}$. However, flows can be neglected ($\boldsymbol{v}=\boldsymbol{0}$) when droplet growth is dominated by diffusion. For example, C. elegans centrosomes and P granules of typical radii $R\approx 1\,\mathrm{\upmu m}$ [109, 110] exhibit velocities of $v\approx 10\,\nicefrac{\mathrm{\upmu m}}{\mathrm{min}}$ [111, 112], but the molecular diffusivity of $D\approx 5\,\nicefrac{\mathrm{\upmu m^{2}}}{\mathrm{s}}$ [113, 114] implies small Péclet numbers ($vR/D\approx 0.03$), and so hydrodynamic transport is less relevant than diffusive transport. Consequently, we also neglect the stochastic component of the hydrodynamic flow [115].
The coupled partial differential equations (2.14) for the fields $\{\phi_{i}\}$ typically permit stationary homogeneous solutions. For instance, without chemical reactions ($s_{i}=0$), each homogeneous state $\phi_{i}(\boldsymbol{r})=\bar{\phi}_{i}$ is stationary because of material conservation. Such a state is stable if the associated Hessian
@@
The late-stage dynamics can in principle be obtained by simulating equations (2.14), which requires two boundary conditions at the system’s boundary because of the fourth-order spatial derivatives. Before we introduce such boundary conditions in section 4, we first consider the simpler case of isolated droplets, where we can employ periodic boundary conditions on all fields. This situation is typically considered in numerical investigations, e.g., using finite differences [116], spectral methods [117, 118], Lattice-Boltzmann methods [119, 120, 121, 122], or stochastic solvers [123, 124]. Since these numerical simulations are costly [125], contemporary computers are not able to simulate multi-component dynamics for the time and length scales relevant in cells [116]. However, to discuss dynamics of individual droplets, it often suffices to consider binary phase separation and sharp interfaces (see section 2.2), which will be the main focus of this review.
@@ 2.1.3  Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black lines) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) for two values of the Laplace pressure $P_{\gamma}$. The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium volume fractions (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
In many cases, we do not need to describe all the individual components of a mixture and can instead focus on droplet material separating from a solvent. In this case, the mixture is treated as a binary fluid, and only the volume fraction $\phi$ of one component needs to be specified (${{N_{\mathrm{c}}}=1}$). The local free energy density then reduces to
@@
and determines the phase behaviour of the liquid; see figure 2.1. In particular, a homogeneous system with volume fraction $\phi$ is stable with respect to small perturbations when the Hessian given in equation (2.15) is negative, i.e., if ${f^{\prime\prime}(\phi)\propto\bar{\mu}^{\prime}(\phi)>0}$. This occurs when the interaction parameter $\chi$ is smaller than
@@
Homogeneous systems are stable against small fluctuations outside the spinodal region, but they are not necessarily the only stable states or even the state with the lowest free energy. To see whether heterogeneous states (e.g., droplets) are able to form, we next consider the equilibrium conditions (2.5). These conditions can be solved graphically using the common–tangent construction (figure 2.1A), or equivalently the equal–area Maxwell construction (figure 2.1B). Alternatively, the equilibrium point can be read off the $\Pi(\bar{\mu})$ plot shown in figure 2.1D, and the corresponding volume fractions then identified via figures 2.1B–C. For the chemical potential of the binary system given by equation (2.17), these graphical constructions are possible, and the solutions unique, if $\chi>\chi_{*}$. The two fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ correspond to the equilibrium fractions of the dilute and dense phase, respectively outside and inside droplets. Note that $\phi^{(0)}_{\mathrm{out}}$ is also known as the saturation concentration since droplets can only form when it is exceeded. The common–tangent construction also illustrates that the phase-separated state generally has lower free energy than the corresponding homogeneous state with equal average composition $\bar{\phi}$. The two states only have equal free energy in the limit where one phase of the phase-separated state vanishes ($\bar{\phi}=\phi^{(0)}_{\mathrm{out}}$ or $\bar{\phi}=\phi^{(0)}_{\mathrm{in}}$). Taken together, the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ as a function of $\chi$ form the binodal line. Since the phase-separated state does not exist outside the binodal line, this line also bounds the region where the homogeneous phase is globally stable; see figure 2.1. The binodal line can be obtained analytically for the symmetric case,
@@
for $\Omega=1$. In the limit of strong interactions ($\chi\gg 1$), this predicts $\phi^{(0)}_{\mathrm{out}}\approx e^{-\chi}$, which reveals that even a minute increase of the interaction energy by $2k_{\mathrm{B}}T$ (so $\chi$ is increased by $2$) leads to an almost $10$-fold decrease in the saturation concentration $\phi^{(0)}_{\mathrm{out}}$. In contrast, the spinodal branch scales as $\phi^{\mathrm{spin}}_{\mathrm{out}}\approx 1/(2\Omega\chi)$, implying that a linear stability analysis of the homogeneous state is inadequate to study phase separation in detail.
The entire phase diagram can be separated into three distinct regions. For low $\chi$ or dilute solutions ($\phi<\phi^{(0)}_{\mathrm{out}}$ or $\phi>\phi^{(0)}_{\mathrm{in}}$) only homogeneous states persist. Conversely, only the phase-separated state exists in the spinodal region when interactions are sufficiently strong ($\chi>\chi_{\mathrm{spin}}$). Between these two regions, homogeneous and phase-separated states are both locally stable, but the phase-separated state has lower energy; nucleation events thus become possible when thermal fluctuations are included.
@@
where we for simplicity neglect hydrodynamic flows $\boldsymbol{v}$, and the mobility $\Lambda(\phi)=\Lambda_{0}\phi(1-\phi)$ follows from equation (2.13). Without the source term $s$, which accounts for reactions (see section 5), equation (2.23) is know as the Cahn–Hilliard equation [99]. This equation describes many phase separation phenomena and has been studied extensively [127].